OpenStudy (anonymous):

Please, would anyone help me with this? I'm looking for a proof that pi is irrational.

OpenStudy (anonymous):

google

OpenStudy (anonymous):

I did already, the proof I saw made a step I can't follow, so I came here to see if anyone ahs a simpler one.

OpenStudy (anonymous):

unlikely. it is not like proving that \(\sqrt{2}\) is irrational. it is going to take some doing no matter which proof you use

OpenStudy (anonymous):

Ok then, so could you help me with this crucial step? "We define the polynomial function: ∀x∈R:f(x)=xn(a−bx)nn! We differentiate this 2n times, and then we build: F(x)=∑j=0n(−1)jf(2j)(x)=f(x)+⋯+(−1)jf(2j)(x)+⋯+(−1)nf(2n)(x) ... that is, the alternating sum of f and its first n even derivatives." I don't get the result of differentiating 2n times, it seems like it's skipping steps.

OpenStudy (anonymous):

the wiki proof i am looking at has something similar, is that what you are looking at? it defines F as the sum of f and its first n even derivatives.

OpenStudy (anonymous):

Yes, I guess it could help.

OpenStudy (anonymous):

http://www.coolissues.com/mathematics/Pi/pi.htm

OpenStudy (anonymous):

Thanks.